# Explicit Abnormality Extraction for Unsupervised Motion Artifact Reduction in Magnetic Resonance Imaging

### v5

@@ 3.2 Loss Functions
$L_{1}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\left|x-y\right|$ (3)
$L_{SSIM}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\left|1-SSIM(x,y)^{2}\right|$ (4)
$L_{adv}(x,D)=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(D(x)-1)^{2}}$ (5)
@@
$L_{G_{r}\_adv}(\overline{x^{a}},D_{b})=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(D_{b}(\overline{x^{a}})-1)^{2}}$ (7)
@@
$L_{cyc}(x^{a},\overline{x^{a}})=L_{1}(x^{a},\overline{x^{a}})+\lambda_{SSIM}*L_{SSIM}(x^{a},\overline{x^{a}})$ (8)
@@ 4.2 Evaluation Metrics
$SSIM(X,Y)=\frac{(2\mu_{X}\mu_{Y}+C_{1})(2\sigma_{XY}+C_{2})}{(\mu^{2}_{X}+\mu^{2}_{Y}+C_{1})(\sigma^{2}_{X}+\sigma^{2}_{Y}+C_{2})}$ (10)
@@
$PSNR=10\log_{10}{\frac{MaxValue^{2}}{MSE}}$ (11)
$MSE=\frac{1}{mn}\sum_{i=0}^{m-1}\sum_{j=0}^{n-1}[I(i,j)-K(i,j)]^{2}$ (12)

### v6

@@ 3.2 Loss Functions
$L_{1}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\left|x-y\right|,$ (3)
$L_{SSIM}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\left|1-SSIM(x,y)^{2}\right|,$ (4)
$L_{adv}(x,D)=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(D(x)-1)^{2}},$ (5)
@@
$L_{G_{r}\_adv}(\overline{x^{a}},D_{b})=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(D_{b}(\overline{x^{a}})-1)^{2}}.$ (7)
@@
$L_{cyc}(x^{a},\overline{x^{a}})=L_{1}(x^{a},\overline{x^{a}})+\lambda_{SSIM}*L_{SSIM}(x^{a},\overline{x^{a}}),$ (8)
@@ 4.2 Evaluation Metrics
$SSIM(X,Y)=\frac{(2\mu_{X}\mu_{Y}+C_{1})(2\sigma_{XY}+C_{2})}{(\mu^{2}_{X}+\mu^{2}_{Y}+C_{1})(\sigma^{2}_{X}+\sigma^{2}_{Y}+C_{2})},$ (10)
@@
$PSNR=10\log_{10}{\frac{MaxValue^{2}}{MSE}},$ (11)
$MSE=\frac{1}{mn}\sum_{i=0}^{m-1}\sum_{j=0}^{n-1}[I(i,j)-K(i,j)]^{2},$ (12)